OpenStudy (vera_ewing):

Which ratio would be used to determine how many moles of water are produced by reacting 15 mol of oxygen in excess hydrogen? 2H2 + O2 2H2O

OpenStudy (vera_ewing):

@JFraser

OpenStudy (jfraser):

if you're asked about the ratio between OXYGEN and the WATER produced, what are the coefficients between those two?

OpenStudy (vera_ewing):

The answer is 2 mol H2O/1 mol O2 right?

OpenStudy (vera_ewing):

Or is it the other way around?

OpenStudy (jfraser):

one mole of oxygen will produce 2 moles of water

OpenStudy (vera_ewing):

Ohh so it's \[\frac{ 1 mol H2O }{ 2 mol O2 }\]

OpenStudy (jfraser):

yes

OpenStudy (vera_ewing):

Wait that's not one of my answer choices though....

OpenStudy (jfraser):

sorry, you flipped the numbers

OpenStudy (jfraser):

it should be \[\frac{1mol O_2}{2mol H_2O}\]

OpenStudy (vera_ewing):

No that's not one of my answer choices either...

OpenStudy (jfraser):

then flip the whole thing, and its still the same ratio\[\frac{2mol H_2O}{1mol O_2}\]

OpenStudy (vera_ewing):

\[\frac{ 2 mol H2O }{ 1 mol O2 }\]

OpenStudy (vera_ewing):

oh ok thanks!

OpenStudy (jfraser):

a ratio that's "upside down" is still the same ratio

OpenStudy (vera_ewing):

ohhh ok thanks for explaining that!